To calculate the number of years until maturity, assume that it is currently January 15, 2013.
Company
(Ticker)
Coupon Maturity Last
Price
Last
Yield
EST $ Vol
(000’s)
  Xenon, Inc. (XIC) 6.600    Jan 15, 2032      94.303        ??        57,374      
  Kenny Corp. (KCC) 7.240    Jan 15, 2029      ??           6.26       48,953      
  Williams Co. (WICO) ??      Jan 15, 2038      94.855        7.08       43,814      
Requirement 1:

What price would you expect to pay for the Kenny Corp. bond? (Do not round intermediate calculations.Round your answer to 2 decimal places (e.g., 32.16).)

  Price $   
Requirement 2:

What is the bond’s current yield? (Round your answer to 2 decimal places (e.g., 32.16).)

  Current yield %
Answer #1

Face Value of kenny Bond = $1000

Annual Coupon Payment = $1000*7.240%

= $72.40

No of Years to Maturity(n) = 16 years

Yield = 6.26%

Calculating the Price of Bond:-

+ ... Coupon Payment (1+YTM) Price Coupon Payment Coupon Payment + (1+YTM) (1 + YTM2 FaceValue (1+YTM)

Price = \frac{72.40}{(1+0.0626)^{1}}+\frac{72.40}{(1+0.0626)^{2}}+.....+\frac{72.40}{(1+0.0626)^{16}}+\frac{1000}{(1+0.0626)^{16}}

Price = $718.78 + $378.514

Price = $1097.29

So, The price of bond of Kenny Corp is $1097.29

b). Current Yield = Coupon payment/Current price

= $72.40/$1097.29

= 6.60%

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating